OpenStudy (anonymous):

How do you solve each system by elimination.? -6x-6y-2z=-12 -6x+y--4z=18 -3x-2y+2z=18

OpenStudy (anonymous):

Okay start off simple so since the first two have the same X I would start with them. -6x - 6y - 2z = -12 (for the second equation just change each sign to the opposite of what it was) 6x - y + 4z = -18 Just add them together to get -> -7y + 2z = -30 Can you do the rest now that you understand elimination?